# Using LLMs in Generating Design Rationale for Software Architecture Decisions

### v1

@@ 5.2. Strengths and Limitations of LLM-Generated DR
The proportion of “Uncertain” arguments is 4.87%, 4.57%, and 4.12% with the three prompting strategies, respectively. Compared to the zero-shot method, neither the CoT method nor LLM-based agents significantly reduce the proportion of “Uncertain” arguments. A potential reason is the insufficient contextual information related to the architecture. Although we provided as much detailed contextual information as possible collected from the data extraction phase, the available information for LLMs to generate DR remains limited. For instance, regarding architecture problems in GitHub projects, LLMs lack critical insights that project developers possess, such as the previous versions and specific code contexts. Moreover, the challenges that architecture design faces in real-world scenarios remain unknown to LLMs. As a result, the “Uncertain” arguments proposed by LLMs require validation in practical production environments. However, “Uncertain” arguments may also offer valuable insights, encouraging architects to further analyze and consider them in their projects context.
@@ 6. Threats to Validity
In this section, we discussed the threats to validity following the guidelines by Runeson and Höst (Runeson and Höst, 2009). It should be mentioned that internal validity is excluded as we did not explore the causal relationships between variables and outcomes.
@@
Construct Validity. The main threat to the construct validity of this study is the evaluation methods we employed to assess the quality of LLM-generated DR. Although we selected architecture decisions with DR provided by human experts in our dataset, it is important to note that there is typically no universally definitive DR for a given architecture decision. Therefore, we employed two complementary methods for evaluating the generated DR: accuracy metrics and the IHUM-category. For accuracy metrics, we compared the arguments in the generated DR with those provided by human experts to calculate the Precison, Recall and F1-score value. Since the DR provided by human experts is used to explain architecture decisions, the values of the accuracy metrics can reflect the capability of LLMs to recover the DR for architecture decisions. However, accuracy metrics considered the DR provided by human experts as the ground truth, overlooking the arguments in LLM-generated DR that may differ from those of human experts. Therefore, we proposed the IHUM-category to qualitatively assess the arguments in the generated DR, categorizing them into those that align with the human experts’ arguments (“Insightful”), those that differ but offer valid perspectives (“Helpful”), those that are ambiguous (“Uncertain”), and those contain erroneous information (“Misleading”). We evaluated the quality of LLM-generated DR by using both methods simultaneously, aiming to assess generated DR as comprehensively as possible.
@@
External Validity. The primary threat to the external validity is the selection of data sources. In our dataset, we chose the posts from SO, as well as the issues and discussions from GitHub to maximize the diversity of data sources. SO is a popular Q&A community where developers engage in discussions, including a wide range of architecture-related issues (de Dieu et al., 2023). GitHub Issues are used for tracking bugs, feature requests, and raising potential issues, while GitHub Discussions facilitate project-related discussions and knowledge sharing. Both the two sources on GitHub can provide architecture decisions from practical projects. Therefore, the three data sources ensure, to a certain extent, the diversity of our dataset. However, we acknowledge that our dataset is not comprehensive and does not cover all architecture problems. Another threat to the external validity is the selection of LLMs. We chose five widely used open-source and closed-source LLMs that have demonstrated strong performance in other SE research. However, the performance of these five LLMs may not fully represent the state-of-the-art capabilities of LLMs in generating DR, as other models with potentially superior performance may not be included. Therefore, due to the ever-evolving nature of LLMs, we plan to continue exploring their potential for generating DR in the future.
@@
Reliability. Since the dataset construction and results evaluation process were conducted manually, there are potential to introduce personal bias. To reduce this threat, the first and fourth authors conducted pilot experiments before data labeling, extraction, and IHUM-category classification. During the data labeling and evaluation of the results (i.e., using accuracy metrics and IHUM-category), the consistency between the first and fourth authors was evaluated using the Cohen’s Kappa coefficient, yielding values of 0.834, 0.863, and 0.738, respectively. The Cohen’s Kappa value indicates an almost perfect agreement between the two authors on data labeling and accuracy metrics, and substantial agreement on the IHUM-category classification results (Landis and Koch, 1977). If any disagreement arose during the pilot and formal process, the second and third author were involved in reaching a consensus. The constructed dataset and evaluation results were reviewed multiple times by the four authors (the first to fourth authors) to ensure that they aligned with the pre-established criteria. Furthermore, we have made available the dataset of the study (Zhou et al., 2025) to enable other researchers to replicate and validate our findings.
@@ 7.1. Design Rationale in Software Architecture
Several studies focused on DR in software architecture. Van der Ven et al. (van der Ven et al., 2006) indicated that explicitly modeling design decisions in software architecture can bridge the gap between rationale management and architecture artifacts, as it enables a close integration of rationale management with the architecture. Tang et al. (Tang et al., 2006) conducted a survey targeting practitioners to explore their perceptions of the value of DR and to understand how they utilize and document the background knowledge associated with their architecture decisions. In another study, Tang et al. (Tang et al., 2007) introduced a rationale-based architecture model that integrates design rationale, design objects, and their relationships to address the issue of design rationale often being undocumented or unstructured. Davide et al. (Falessi et al., 2008) proposed a new approach to Design Decision Rationale Documentation (DDRD), which tailors the documentation based on its intended use or purpose. In another study, Davide et al. (Falessi et al., 2013) conducted an empirical study through two controlled experiments and proved that the value of a Design Rationale Documentation (DRD) information item depends on its category (e.g., Assumptions, Related Requirements) and the activity it supports. Soliman et al. (Soliman, 2024) analyzed 156 architectural emails to identify the design rationale used in mailing lists and their relationship with decision types. They identified nine types of design rationale, six relationships between rationale types, and three relationships between decision and rationale types. Zhao et al. (Zhao et al., 2024) developed DRMiner, to automatically mine latent design rational from the discussion of developers in open-source community.
@@ 7.2. LLM-based Understanding of Software Artifacts
Numerous studies have been conducted to explore the capability of LLMs to understand various software artifacts. Xie et al. (Xie et al., 2024) evaluated the ability of LLMs to generate software specifications from software comments or documentation using few-shot learning. They compared the performance of 13 state-of-the-art LLMs with traditional methods across three publicly available datasets. Additionally, they conducted a comparative analysis of the failure cases from LLMs and traditional approaches, highlighting their respective strengths and weaknesses. Nam et al. (Nam et al., 2024) developed an LLM-based conversational user interface to assist users in understanding code. The results show that using their plugin is more effective in code completion than relying on web searches. Shaike et al. (Shaik et al., 2024) presented S3LLM, a framework based on LLMs, to understand large-scale scientific software using source code , code metadata, and summarized information from textual technical reports. Pex (Pex, 2024) studied the application of LLMs in software explanation and proposed the FLASE tool, which combines RAG and knowledge graphs, to enhance developers’ understanding of software systems. Pan et al. (Pan et al., 2025) conducted an empirical study to evaluate the capability of LLMs to understand design patterns. Franciscatto Guerra et al. (Guerra and Ernst, 2025) investigated the ability of an LLM to comprehend, replicate, and create structures within the intricate VIPER architecture, a design pattern used in iOS application development. Their results emphasize the potential of LLMs to lower development costs, as well as the challenges the LLMs face in being effectively applied to real-world software design situations. Casillo et al. (Casillo et al., 2025) constructed a dataset of 45,945 commits, each accompanied by rationale explaining the code changes. They then trained a model for generating rationale of code changes, highlighting the challenges in automating this process. Soliman et al. (Soliman and Keim, 2025) evaluated LLM’s understanding of software architecture by comparing its responses to a predefined ground truth. They found that while GPT offers initial insights, expert validation is still needed for reliable results of architecture knowledge. In their study, LLMs were also employed to generate DR for architecture decisions in HDFS.
@@ 7.3. Conclusive Summary
Previous studies (van der Ven et al., 2006; Tang et al., 2006, 2007) have highlighted the significant value of DR in software architecture. Besides, the application of LLM-based methods for understanding software artifacts (Xie et al., 2024; Pex, 2024; Soliman and Keim, 2025) has demonstrated the potential of LLMs in comprehending software architecture and generating DR. However, there has been limited focus on examining the capability of LLMs to generate DR for architectural decisions. Although the study of Soliman et al. (Soliman and Keim, 2025) covered querying LLMs about DR related to software architecture of HDFS, they mainly focused on examining whether LLMs possess specific architectural knowledge using the zero-shot method, rather than on the ability of LLMs to recover DR according to architectural problems and decisions. In our research, we placed greater emphasis on examining the capability of LLMs to generate DR for diverse real-world architecture problems and corresponding architecture decisions by collecting data from three sources (i.e., SO, GitHub Discussions, and GitHub Issues). Furthermore, we employed two complementary methods (accuracy metrics and the IHUM-category) to evaluate how the optimization of prompting strategies can enhance the quality of the generated DR by employing three different prompting strategies (i.e., zero-shot, CoT, and LLM-based agents).
@@ 8. Conclusions and Future Work
With the evolution of LLMs, we plan to continuously explore the performance of emerging LLMs in DR generation. Additionally, according to the limitations observed in the five selected LLMs for DR generation, we aim to further optimize our LLM-based agent system. To be specific, we intend to enhance Information_Collector agent with the capability to locate code contexts relevant to architecture decisions within software projects, thereby providing Aspect_Analyst agent with specific code contexts for analysis. We also aim to augment Aspect_Analyst agent with the ability to identify the external knowledge underlying the generated arguments, as well as automatically generate code examples based on the arguments, to help LLM-based agents provide DR closer to practical development scenarios.
@@
The replication package for this work has been made available at (Zhou et al., 2025).

### v3

@@ 5.2. Strengths and Limitations of LLM-Generated DR
The proportion of “Uncertain” arguments is 4.87%, 4.57%, and 4.12% with the three prompting strategies, respectively. Compared to the zero-shot method, neither the CoT method nor LLM-based agents reduce the proportion of “Uncertain” arguments. A potential reason is the insufficient contextual information related to the architecture. Although we provided as much detailed contextual information as possible collected from the data extraction phase, the available information for LLMs to generate DR remains limited. For instance, regarding architecture problems in GitHub projects, LLMs lack critical insights that project developers possess, such as the previous versions and specific code contexts. Moreover, the challenges that architecture design faces in real-world scenarios remain unknown to LLMs. As a result, the “Uncertain” arguments proposed by LLMs require validation in practical production environments. However, “Uncertain” arguments may also offer valuable insights, encouraging architects to further analyze and consider them in their project context.
5.3. Implications from Interviews
Implication 1: Practitioners consider LLMs as an efficient assistant in generating and recovering DR for architecture decisions, rather than a fully trustworthy expert.
By inquiring into the trustworthiness and applicability of LLM-generated DR, we found that practitioners regarded DR as a useful reference and were able to identify appropriate scenarios for practical application. However, practitioners typically regarded LLM-generated DR as a supplementary aid to support their architectural understanding, rather than as authoritative expert advice. Specifically, half of the practitioners in our interviews stated that trusting LLM-generated DR requires substantial human review. Other practitioners expressed a higher trust level, but none would “fully trust it without human validation”. From the perspective of practitioners, LLM-generated DR is seen as less trustworthy than that provided by experienced human experts. For this reason, P5 noted that he was “not sure” whether he would rely on LLMs’ DR when working within a domain his team is already familiar with. To enhance practitioners’ trust in LLM-generated DR, two key challenges need to be addressed. First, it is essential to demonstrate that LLMs are capable of understanding complex project-related contexts - one of the primary concerns raised by practitioners during the interviews. Without such capability, DR generated by LLMs may be perceived as simplistic, thereby limiting their value in supporting practitioners’ in-depth understanding of architecture design. Second, building trust in LLM-generated DR requires extensive empirical validation in practical software development settings. Practitioners emphasized the importance of evaluating how LLM-generated DR performs in real development processes, particularly in terms of accurately revealing the rationale of architecture decisions and supporting meaningful reasoning across diverse project contexts.
Implication 2: Providing LLMs with project-specific context is highly beneficial, yet challenging to achieve due to the associated costs.
During our interviews, all participants consistently emphasized that providing LLMs with detailed project context significantly enhances the quality and usefulness of the generated DR. However, they also pointed out that the cost of collecting and providing such contextual information is non-negligible for architects, which is also the most frequently mentioned challenge across the interviews. For example, P5 noted that “if the process of collecting project context is too burdensome, it could outweigh the benefits of using LLMs to generate DR”. To address this issue, a potential solution is to further enhance the capability of Information_Collector agent to extract project-specific context. In this study, Information_Collector agent was designed to dynamically retrieve contextual information using a search engine, in order to support diverse architectural scenarios for different projects. However, in real-world software development, practitioners typically focus on understanding the architecture of a single project. Therefore, it is feasible to customize Information_Collector agent for a specific project by targeting the artifacts of the project (e.g., code, documentation, and configuration files). By aligning the extraction process with the architect’s specific information needs, Information_Collector agent can automatically retrieve key architecture-relevant context. This capability is crucial for improving both the quality and practical utility of LLM-generated DR.
Implication 3: Enhancing LLM-based agents with visual output capabilities can facilitate comprehension of DR.
In our interviews, P1 and P4 commented on the textual form of DR generated by LLMs. P1 noted that, compared to DR provided by human experts, the longer textual outputs generated by LLMs imposed a certain cognitive load, which made him “somewhat reluctant to read them in full”. In addition, P5 argued that the visual form may be more effective than textual explanations in some cases, which enables architects to “quickly comprehend the architecture and the rationale behind it”. Similarly, according to the experiment conducted by Shahin et al. (shashin2014do), visualization of architectural design decisions and their DR can improve the understanding of architecture design. Consequently, a promising direction for our future work lies in equipping LLM-based agents with the ability to generate visual representations of architecture decisions with their corresponding DR, which can better assist architects in comprehending DR and mitigating cognitive burden.
@@ 6. Threats to Validity
In this section, we discussed the threats to validity following the guidelines by Runeson and Höst (runeson2009guidelines). It should be mentioned that internal validity is excluded, as we did not explore the causal relationships between variables and outcomes.
@@
Construct Validity. The main threat to the construct validity of this study is the evaluation methods we employed to assess the quality of LLM-generated DR. Although we selected architecture decisions with DR provided by human experts in our dataset, it is important to note that there is typically no universally definitive DR for a given architecture decision. Therefore, we employed two complementary methods for evaluating the generated DR: accuracy metrics and the IHUM-category. For accuracy metrics, we compared the arguments in the generated DR with those provided by human experts to calculate the Precision, Recall, and F1-score value. Since the DR provided by human experts is used to explain architecture decisions, the values of the accuracy metrics can reflect the capability of LLMs to recover the DR for architecture decisions. However, accuracy metrics considered the DR provided by human experts as the ground truth, overlooking the arguments in LLM-generated DR that may differ from those of human experts. Therefore, we proposed the IHUM-category to qualitatively assess the arguments in the generated DR, categorizing them into those that align with the human experts’ arguments (“Insightful”), those that differ but offer valid perspectives (“Helpful”), those that are ambiguous (“Uncertain”), and those contain erroneous information (“Misleading”). We evaluated the quality of LLM-generated DR by using both methods simultaneously, aiming to assess generated DR as comprehensively as possible. Another potential threat to construct validity arises from Information_Collector agent, which gathers relevant knowledge via search engines. Although we excluded URLs that overlapped with those in our dataset to avoid direct access to original data entries, content mirroring on other websites (e.g., from SO) may still result in unintended information leakage. To fully avoid this issue, search engine date-range restrictions can be used to ensure all retrieved information predates the creation of the architecture problem. However, this approach may also limit the external knowledge Information_Collector agent can obtain from search engines, as some potentially useful recent information might be filtered out. In our future work, we can optimize the filtering strategy to effectively minimize information leakage while maximizing the capture of potentially valuable external knowledge for Information_Collector agent.
@@
External Validity. The primary threat to the external validity is the selection of data sources. In our dataset, we chose the posts from SO, as well as the issues and discussions from GitHub to maximize the diversity of data sources. SO is a popular Q&A community where developers engage in discussions, including a wide range of architecture-related issues (musengamana2023characterizing). GitHub Issues are used for tracking bugs, feature requests, and raising potential issues, while GitHub Discussions facilitate project-related discussions and knowledge sharing. Both the two sources on GitHub can provide architecture decisions from practical projects. Therefore, the three data sources ensure, to a certain extent, the diversity of our dataset. However, we acknowledge that our dataset is not comprehensive and does not cover all architecture problems. Another threat to the external validity is the selection of LLMs. We chose five widely used open-source and closed-source LLMs that have demonstrated strong performance in other SE research. However, the performance of these five LLMs may not fully represent the state-of-the-art capabilities of LLMs in generating DR, as other models with potentially superior performance may not be included. Moreover, based on the results of five selected LLMs, it remains difficult to reveal the relationship between the quality of generated DR and model-specific attributes, such as parameter size, model architecture, or release date. On the one hand, the outputs of LLMs are inherently difficult to interpret; on the other hand, the number of models included in our experiments is limited, making it difficult to perform a generalizable analysis. For example, gpt-4-0613 achieves the best performance in terms of Recall, as well as in generating “Helpful” arguments while avoiding “Uncertain” and “Misleading” ones. However, due to the lack of access to gpt-4-0613’s internal mechanisms and training data, as well as the absence of comparative experiments (e.g., control for model size), it is challenging to rigorously explain the reasons behind its superior performance. In addition, as discussed in Section 3.2, there is a trade-off between ensuring a sufficiently large dataset and selecting the most recent LLMs. As a result, the five selected LLMs in our study were released over a span of one year and one month. Although our experimental results did not show an advantage of the more recently released model (i.e., llama3-8B) over others, the inconsistency in the release periods of the LLMs remains a potential threat to external validity. Therefore, we plan to continue exploring the capabilities of a broader range of models released across different time periods for generating DR in our future work, and to design controlled comparative experiments to systematically investigate how model-specific attributes - such as size, architecture, and training data - influence DR quality.
@@
Reliability. Since the dataset construction and results evaluation process were conducted manually, there are potential to introduce personal bias. To reduce this threat, the first and fourth authors conducted pilot experiments before data labeling, extraction, and IHUM-category classification. During the data labeling and evaluation of the results (i.e., using accuracy metrics and IHUM-category), the consistency between the first and fourth authors was evaluated using the Cohen’s Kappa coefficient, yielding values of 0.834, 0.863, and 0.738, respectively. The Cohen’s Kappa value indicates an almost perfect agreement between the two authors on data labeling and accuracy metrics, and substantial agreement on the IHUM-category classification results (landis1977theMO). If any disagreement arose during the pilot and formal process, the second and third authors were involved in reaching a consensus. The constructed dataset and evaluation results were reviewed multiple times by the four authors (the first to fourth authors) to ensure that they aligned with the pre-established criteria. However, since both data labeling and evaluation were conducted by the authors, the Kappa values - although indicative of almost perfect agreement between the first and fourth authors - may not fully guarantee objectivity. While the Kappa statistic measures the degree of agreement among annotators, it does not account for potential systematic biases. For instance, if all annotators exhibit similar interpretive tendencies, the high agreement may reflect shared subjective perspectives rather than genuine objectivity. To reduce such a threat to reliability and foster reproducibility, we have provided the pilot and formal results of data labeling and evaluation in the dataset of this study (dataset), including the pilot and formal results of data labeling and evaluation, enabling other researchers to replicate and validate our findings.
@@ 7.1. Design Rationale in Software Architecture
Several studies focused on DR in software architecture. Van der Ven et al. (vanderVen2006design) indicated that explicitly modeling design decisions in software architecture can bridge the gap between rationale management and architecture artifacts, as it enables a close integration of rationale management with the architecture. Tang et al. (tang2006asurvey) conducted a survey targeting practitioners to explore their perceptions of the value of DR and to understand how they utilize and document the background knowledge associated with their architecture decisions. In another study, Tang et al. (tang2007rationale) introduced a rationale-based architecture model that integrates design rationale, design objects, and their relationships to address the issue of design rationale often being undocumented or unstructured. Davide et al. (davide2008value) proposed a new approach to Design Decision Rationale Documentation (DDRD), which tailors the documentation based on its intended use or purpose. In another study, Davide et al. (davide2013value) conducted an empirical study through two controlled experiments and proved that the value of a Design Rationale Documentation (DRD) information item depends on its category (e.g., Assumptions, Related Requirements) and the activity it supports. Soliman et al. (soliman2024exploring) analyzed 156 architectural emails to identify the design rationale used in mailing lists and their relationship with decision types. They identified nine types of design rationale, six relationships between rationale types, and three relationships between decision and rationale types. Zhao et al. (zhao2024drminer) developed DRMiner, to automatically mine latent design rationale from the discussion of developers in open-source communities. Meanwhile, design rationale is a core element of architectural design decisions (shahin2009architectural). Dhar et al. (dhar2024can) explored the feasibility of using LLMs to generate architectural design decisions based on contextual information, providing preliminary evidence of LLMs’ ability to assist in generating design rationale. Building on this line of work, Dhar et al. (dhar2025drafting) further investigated few-shot prompting, retrieval-augmented generation, and fine-tuning to enhance LLM-based generation of architecture decisions, and proposed DRAFT, which outperforms prior approaches in generating architecture decisions on a dataset of 4,911 Architecture Decision Records (ADRs). Díaz-Pace et al. (díaz-pace2024helping) proposed ArchMind, a generative AI-based design assistant that leverages architectural knowledge and system information to help novice architects select and assess alternative architecture decisions, generating well-justified ADRs in a classroom experiment.
@@ 7.2. LLM-based Understanding of Software Artifacts
Numerous studies have been conducted to explore the capability of LLMs to understand various software artifacts. Xie et al. (xie2024leveraging) evaluated the ability of LLMs to generate software specifications from software comments or documentation using few-shot learning. They compared the performance of 13 state-of-the-art LLMs with traditional methods across three publicly available datasets. Additionally, they conducted a comparative analysis of the failure cases from LLMs and traditional approaches, highlighting their respective strengths and weaknesses. Nam et al. (nam2024using) developed an LLM-based conversational user interface to assist users in understanding code. The results show that using their plugin is more effective in code completion than relying on web searches. Shaike et al. (shaik2024s3s3llm) presented S3LLM, a framework based on LLMs, to understand large-scale scientific software using source code, code metadata, and summarized information from textual technical reports. Pex (chaudron2024exploring) studied the application of LLMs in software explanation and proposed the FLASE tool, which combines RAG and knowledge graphs, to enhance developers’ understanding of software systems. Pan et al. (pan2025codellmsunderstanddesign) conducted an empirical study to evaluate the capability of LLMs to understand design patterns. Franciscatto Guerra et al. (guerra2025acccessing) investigated the ability of an LLM to comprehend, replicate, and create structures within the intricate VIPER architecture, a design pattern used in iOS application development. Their results emphasize the potential of LLMs to lower development costs, as well as the challenges the LLMs face in being effectively applied to real-world software design situations. Casillo et al. (casillo2025towards) constructed a dataset of 45,945 commits, each accompanied by a rationale explaining the code changes. They then trained a model for generating a rationale for code changes, highlighting the challenges in automating this process. Soliman et al. (soliman2025do) evaluated LLMs’ understanding of software architecture by comparing their responses to a predefined ground truth. They found that while GPT offers initial insights, expert validation is still needed for reliable results of architecture knowledge. In their study, LLMs were also employed to generate DR for architecture decisions in HDFS.
@@ 7.3. Conclusive Summary
Previous studies (vanderVen2006design; tang2006asurvey; tang2007rationale) have highlighted the significant value of DR in software architecture. Besides, the application of LLM-based methods for understanding software artifacts (xie2024leveraging; chaudron2024exploring; soliman2025do) has demonstrated the potential of LLMs in comprehending software architecture and generating DR. However, there has been limited focus on examining the capability of LLMs to generate DR for architecture decisions. Although the study of Soliman et al. (soliman2025do) covered querying LLMs about DR related to the software architecture of HDFS, they mainly focused on examining whether LLMs possess specific architectural knowledge using the zero-shot method, rather than on the ability of LLMs to recover DR according to architecture problems and decisions. Dhar et al. (dhar2024can; dhar2025drafting) concentrated on improving the quality of architecture decisions generated by LLMs through various prompting strategies; however, their study did not provide a systematic examination of the quality of the corresponding DR. Díaz-Pace et al. (díaz-pace2024helping) proposed a generative AI–based design assistant (ArchMind) aimed at supporting novice architects. While their results suggest that LLMs can produce architecture decisions with reasonable rationale, the study did not extend to a systematic analysis of the generated DR or to an explicit assessment of LLMs’ ability to recover or generate DR. In our research, we placed greater emphasis on examining the capability of LLMs to generate DR for diverse real-world architecture problems and corresponding architecture decisions by collecting data from three sources (i.e., SO, GitHub Discussion, and GitHub Issues). Furthermore, we employed two complementary methods (accuracy metrics and the IHUM-category) to evaluate how the optimization of prompting strategies can enhance the quality of the generated DR by employing three different prompting strategies (i.e., zero-shot, CoT, and LLM-based agents).
@@ 8. Conclusions and Future Work
Based on our semi-structured interviews, practitioners found LLM-generated DR valuable, but some noted that it might require substantial human review to be fully trusted. They emphasized that providing project context, architectures and DR from similar projects, and domain-specific knowledge could enhance the credibility of the generated DR. Most practitioners expressed willingness to adopt LLMs for DR generation in practice, and the participants identified four typical application scenarios for applying generated DR: comparing alternative architecture designs, learning unfamiliar architectures, gaining deeper insights into current architecture decisions, and guiding architectural evolution. At the same time, they highlighted four major challenges, which include the difficulty of providing LLMs with sufficient project context, limited understanding of complex requirements by LLMs, concerns about the reliability of LLM-generated DR, and the restriction of LLM’s outputs in textual form.
With the evolution of LLMs, we plan to continuously explore the performance of emerging LLMs in DR generation. Additionally, according to the limitations observed in the five selected LLMs for DR generation, we aim to further optimize our LLM-based agent system. To be specific, we intend to enhance Information_Collector agent to automatically gather project-relevant context, including code, documentation, and configuration files, in order to reduce the burden on practitioners while providing Aspect_Analyst agent with the necessary project context for generating DR. We also aim to augment Aspect_Analyst agent with the ability to identify the external knowledge underlying the generated arguments, as well as automatically generate code examples based on the arguments, to help LLM-based agents provide DR closer to practical development scenarios. Moreover, our future work will explore equipping LLM-based agents with the ability to generate visual representations of DR, which may facilitate comprehension and alleviate cognitive load for practitioners.
@@
The replication package for this work has been made available at (dataset).